# Integrating Open-World Shared Control in Immersive Avatars

### v1

@@ I Introduction
Teleoperation allows humans to sense and act in remote locations that may be hazardous or difficult to access. Recently, several groups have developed robot avatars [29, 22, 20, 4] that provide immersive interfaces for operators to control an entire robot body and transport their presence to a remote location. These systems have proven that avatars enable novice operators to intuitively inspect, navigate, and manipulate the remote environment, but even state-of-the-art systems lag behind human proficiency [10].
@@
Figure 1: Avatar robot unscrewing a jar (a) while controlled by an operator (b). The operator’s view is overlaid with a predictive menu (c) that suggests possible assistive actions and shows corresponding affordances as augmented-reality objects (purple circle overlaying the jar lid). The operator can use a laser pointer attached to their controller to select one of the suggested actions (d). [Best viewed in color.]
@@
This skill gap has long been identified as an issue for teleoperation, and researchers have proposed many assistance schemes to mitigate it, including virtual fixtures [26, 1, 24, 12], mode switches [25], and automated planning [17, 18, 5]. Assistance has been shown to help operators in structured lab settings, but several challenges remain before they can be deployed, such as “open-world” tasks [32], predicting the operator’s intent [19], evaluating and managing the operator’s trust [19], and operator overload degrading the operator’s fluency [7]. The open-world problem is particularly troublesome, since teleoperation is especially effective in leveraging human problem-solving and contextual understanding, but nearly all assistance methods are designed to work with predefined objects in semi-structured scenarios [5, 25, 12, 31]. Another major challenge is bridging assistance paradigms with the immersive paradigm. Existing avatars incorporate few assistive features [28, 20, 21], whereas shared control literature typically considers non-immersive mouse and keyboard interfaces [18, 24]. The question of how to integrate these schemes introduces several design challenges, such as how to allow the operator to quickly switch between control modes and to configure different types of assistance without occluding the view of the remote environment.
@@
The goal of this work is to design and evaluate a framework to incorporate open-world shared control into immersive robot avatars. To address the central design challenges highlighted above, we designed an in-headset menu that allows the operator to launch and configure assistive actions using the same interfaces they use to directly control the robot (Fig. 1). Inspired by [24], we implement assistive actions based on geometric affordances that are agnostic to object identity, allowing them to work in a wide range of scenarios. Affordances are rendered as augmented reality (AR) markers in the operator’s immersive view when the user is configuring action targets. We further enhance the fluency of this interface using an “autocomplete” predictive menu paradigm that predicts the operator’s intent in the context of the current scene and history [23]. We incorporate this framework into an avatar system and evaluate novice users on long-form tasks that require many uses of the assistive actions. Human subjects testing $(N=19)$ verifies that our approach, with and without the predictive menu, improves task success rates, completion times, workload, and system usability over standard direct control interfaces while preserving the operator’s sense of presence.
@@ II Related Work
The recent ANA Avatar XPRIZE competition spurred rapid development of teleoperated avatar robots capable of transporting basic human manipulation skills to remote environments [10]. As the competition emphasized immersion and presence, most teams made very little or no use of shared control, instead opting to give as much direct control to the operator as possible. This choice makes the systems open-world, immersive, and intuitive, but users still struggle to perform tasks through the robot as proficiently as they would in-person [10]. Shared control methods could hypothetically assist in operator proficiency while preserving desirable aspects of immersion, but mechanisms for achieving such integration are not well studied.
Operator assistance for non-immersive interfaces has received much attention in the literature. A significant line of work addresses reaching for an object [6], especially when the operator’s interface has fewer DoFs than the robot [11, 13, 25, 14]. In the avatar context, this is not normally a concern because the operator has access to high DoF input devices. Other research provides assistance for complex tasks but requires pre-programmed information about the environment and target objects [25, 5, 12]. For example, [25] presents a system that can perform complicated tasks like opening a door, but key frames of reference for specific objects are labeled by hand, and the state-machines describing transitions between different phases of the tasks are pre-specified. Our work seeks to relax this requirement and provide assistance in an open-world by using more generic types of assistance, detecting affordances at runtime rather than hand labelling them at design-time.
The work of Pruks and Ryu [24] is most similar to our system. Their work uses off-the-shelf methods to segment the environment into geometric primitives and allows the operator to apply customizable virtual fixtures between features detected in the environment and from the robot’s tool. They show that their method improves some task-specific metrics, such as contact force and operator workload, compared to other methods of virtual fixture specification. However, compared to the work here, less attention is paid to the overall robot control interface, requiring the operator to switch between a screen and mouse to specify the virtual fixtures and a haptic device to input low-level motion commands. In contrast, our system allows the operator to use a consistent input interface throughout operation. The tasks used in [24] also did not require operators to modify virtual fixtures once they had been selected, making interface fluency less of a concern. Here, we consider tasks that require frequent changes to the desired virtual fixtures to provide the correct kind of assistance. We also present a framework for incorporating predictive assistance into our system, which Pruks and Ryu did not consider.
@@ III Interface Design
Figure 2: System diagram showing how different interface elements control the robot. Operators use their own head and hand to control the robot’s head and hand, and use a button on their hand controller to interact with the assistive menu. [Best viewed in color.]
@@
We build our work on the TRINA avatar system [21], in which the robot is comprised of two Franka Emika Panda arms, a Robotiq 2F-140 parallel-jaw gripper, an anthropomorphic Psyonic Ability Hand, a Waypoint Vector omnidirectional wheeled base, and a custom-built three DoF neck and head assembly. A human operator controls TRINA using a virtual reality (VR) head-mounted display (HMD) that shows the view of TRINA’s environment from stereo head cameras. They control the robot’s head directly via HMD motion and use VR controllers to move the arms. The operator station is connected to the Internet via Ethernet and the robot is connected via WiFi or an Ethernet tether.
@@ III-A Direct Teleoperation (DT)
The default control mode is the direct teleoperation scheme described in [21]. To simplify novice operator training, in our experiments, we only activate the robot’s right arm, parallel-jaw gripper, and head. The operator wears a VR HMD and the robot’s head tracks the operator’s head orientation. The operator uses a clutched system to control the arm: while holding down a foot pedal, the operator moves a VR controller, shown in Fig. 2, to move the robot’s hand target. This motion is computed relative to the controller’s pose when the operator first presses the pedal. A lower-level controller then attempts to reach this target. The operator can also velocity-control the parallel-jaw gripper using a joystick on the controller, pushing it right to inch the gripper closed, and left to inch it open.
The robot estimates the force applied to its end effector to provide force feedback via two modalities: First, the controller vibrates with an intensity proportional to the estimated force magnitude (clipped between 10 and 30 N). Second, a virtual red hemisphere around the operator’s controller shows the direction of the applied force, and becomes more opaque as the magnitude of the force increases.
@@ III-B Manual Menu (MM)
Using the direct teleoperation interface alone, operators can achieve some manipulation [21], but complicated tasks, such as writing, are still quite difficult. To aid the operator, we created an interface to allow them to execute assistive actions. Guided by previous research [16], we designed a hierarchical pie menu fixed to the operator’s head, shown in Fig. 3. By making the menu hierarchical, we minimize the number of simultaneously displayed icons to keep the operator’s view of the remote environment unobstructed. The operator interacts with the menu using a “laser pointer” emanating from their controller to point at different icons, and clicks the B button on their controller to select them. The operator can bring up this menu by clicking the B button at any time and can close it by selecting the “Close” icon. This menu design allows the operator to configure the menu using the same interface they use to provide low-level commands to the robot, eliminating any need to switch between interfaces during operation. Clicking other icons gives the operator access to different submenus.
@@ III-C Predictive Menu (PM)
While the manual menu provides access to all possible actions, it can be overwhelming and slow, especially for novice users. To alleviate this, we designed a third interface that uses an action predictor, described in section V, to predict the operator’s intent and present them with a reduced menu that only includes the four most likely actions. If the operator’s desired action is not in this set, they can still access the manual menu as a fallback. With this menu, when the operator clicks B, the top four actions are shown instead of the manual menu, as shown in Fig. 1 and Fig. 3. Whenever the operator hovers over an icon corresponding to an action, all other icons (and affordances) dim to 10% opacity. Selecting any icon closes the menu and sends the action to the robot which then executes it.
@@ IV Assistive Actions
We implemented three kinds of assistive actions: constrained teleoperation, snapping to planes, and snapping to circles. Each action accepts a set of parameters determining its precise operating mode. The constrained teleoperation and plane snapping actions were previously described in [23], and so are only briefly covered here.
@@
The constrained teleoperation action, teleop(sens, x, y, z, roll, pitch, yaw) accepts 7 Boolean parameters modifying the operator’s direct control of the arm. During this action, the operator controls the gripper’s target pose in task space by moving a VR controller with their own arm. When the sens parameter is true, the arm’s end-effector motion is scaled to $0.25$ of the operator’s input motion to enable precise manipulation. The remaining parameters toggle constraints on the end-effector motion, activating guidance virtual fixtures to simplify operation [1].
@@
The plane snapping action, snap_to_plane(p) accepts a plane detected from a point-cloud of the environment by a clustering method [8]. This point-cloud is sensed by the “affordance camera” shown in Fig. 2, an Intel RealSense L515 mounted below the robot’s neck, pointed at the center of the robot’s workspace. This action aligns the forward direction of the gripper with the normal of the detected plane and moves it so that its tool tip is $d_{s}=0.15\text{ cm}$ away from the plane. Fig. 4 illustrates this process in 2D. The robot uses a sampling-based planner to find a path to reach this target or reports that no path was found after 10 s.
@@
Lastly, the circle snapping action, snap_to_circle(c), accepts a circle detected from the environment, aligns the gripper’s forward direction with the circle’s axis, and centers the gripper on the circle. Our system detects circles from RGBD images captured by the affordance camera. The system segments the RGB image using the Segment Anything Model (SAM) [15] and also converts the RGBD image into a point-cloud. For each mask found by SAM, the corresponding points are selected, and within this segment the plane supported by the largest number of points is found. The inliers of this plane are computed as the points in the mask within $d_{\text{inlier}}=5\text{ mm}$ of the plane and projected to the plane. The convex hull of these projected points is found and the circle is discarded if this hull’s “circularity” ($\frac{4\pi\cdot\text{Area}}{\text{Perimeter}^{2}}$ [2]) is below $c_{\text{min}}=0.9$. The minimum enclosing circle of the hull is computed and circles with radii greater than $r_{\text{max}}=7\text{ cm}$ are discarded. To avoid duplicate detections, this candidate circle is compared against previously detected circles. Circles are considered similar if the masks from which they were detected overlap, their centers are within $\Delta_{\text{c}}=5\text{ cm}$, and their radii are within $\Delta_{\text{rad}}=1\text{ cm}$. Among all similar circles, only the one with the largest inlier ratio (ratio of inliers to points in the mask) is kept. Once a circle has been selected, the robot computes a target end-effector pose in the same manner as the snap_to_plane action, additionally moving the target so that the projection of the tool tip to the plane of the circle coincides with the circle’s center. Fig. 4 demonstrates this action in 2D.
@@
Figure 4: 2D illustration of the snap_to_plane and snap_to_circle actions. Both align TRINA’s gripper with the normal of the selected affordance, but snap_to_circle centers the gripper on the circle while snap_to_plane only moves it closer to the plane. [Best viewed in color.]
@@ V Intent Prediction
To populate the predictive menu, we require an action predictor that can predict multiple likely actions. Additionally, since the set of affordances is not known until runtime, the predictor must be open-world, i.e. able to predict over an open set of objects. We employ the structured prediction method of [23] as it was found to have strong performance in open-world scenarios on similar tasks.
Actions are defined by a type and a collection of parameters, $\overline{\psi}$, which may be different for each action type. We limit the set of $n$ types a priori and dynamically detect the set of feasible parameters for each type, corresponding to detected affordances. To predict an action given the robot’s current context, $x$, the method uses $n$ parameter scoring neural networks, $\{G^{(i)}(x,\overline{\psi})\}_{i=1}^{n}$, and an action network, $A(x)$. $A(x)$ produces $n$ outputs with each element representing the overall score for an action type. Each $G^{(i)}(x,\overline{\psi})$ predicts a scalar score for parameter collections of a particular action type. To assign a score to a complete action, the appropriate scores are summed together, $s=e_{i}^{\intercal}A(x)+G^{(i)}(x,\overline{\psi})$, where $e_{i}$ is the $i$th standard basis vector.
To train and evaluate our predictor, three expert operators (paper authors) collected a dataset of 150 action sequences across three different tasks: unscrewing a jar lid, writing “IML” on a whiteboard, and plugging a cord into an electrical socket. Each sequence was collected in a highly cluttered environment that contained many different distractor objects with varied compositions and arrangements. The specific target objects used were also modified (for example, varying which jars were used). The scoring function was trained using a maximum margin loss function to output high scores for actions observed in the demonstrations [23].

### v2

@@ I Introduction
Teleoperation allows humans to sense and act in remote locations that may be hazardous or difficult to access. Recently, several groups have developed robot avatars [schwarz_nimbro_2021, marquescommodity, luo_team_2023, vanbotics] that provide immersive interfaces for operators to control an entire robot body and transport their presence to a remote location. These systems have proven that avatars enable novice operators to intuitively inspect, navigate, and manipulate the remote environment, but even state-of-the-art systems lag behind human proficiency [XPRIZESystemsPaper2023].
(a) Robot holding jar in left gripper
(b) Operator interface
(c) Operator’s view with a predictive menu showing suggested actions
(d) A “laser pointer” paradigm is used to select actions
@@
Figure 1: An operator uses the Avatar robot to unscrew a jar using the immersive interface. The predictive menu suggests possible assistive actions and shows corresponding affordances as augmented-reality objects (purple circle overlaying the jar lid). [Best viewed in color.]
@@
This skill gap has long been identified as an issue for teleoperation, and researchers have proposed many assistance schemes to mitigate it, including virtual fixtures [rosenberg_virtual_1993, bowyerActiveConstraintsVirtual2014, pruks_method_2022, huang_evaluation_2019], mode switches [quereSharedControlTemplates2020], and automated planning [leeperMethodsCollisionfreeArm2013, leeperStrategiesHumanintheloopRobotic2012, bustamante_cats_2022]. Assistance has been shown to help operators in structured lab settings, but several challenges remain before they can be deployed, such as “open-world” tasks (tasks where the number and/or types of objects in the robot’s environment are not known a-priori) [young_review_2020], predicting the operator’s intent [li_classification_2023], evaluating and managing the operator’s trust [li_classification_2023], and operator overload degrading the operator’s fluency [fallonArchitectureOnlineAffordancebased]. The open-world problem is particularly troublesome, since teleoperation is especially effective in leveraging human problem-solving and contextual understanding, but nearly all assistance methods are designed to work with predefined objects in semi-structured scenarios [bustamante_cats_2022, quereSharedControlTemplates2020, huang_evaluation_2019, wangTaskAutocorrectionImmersive]. Another major challenge is bridging assistance paradigms with the immersive paradigm. Existing avatars incorporate few assistive features [schwarz_robust_2023, luo_team_2023, AVATRINASystemsPaper], whereas shared control literature typically considers non-immersive mouse and keyboard interfaces [leeperStrategiesHumanintheloopRobotic2012, pruks_method_2022]. The question of how to integrate these schemes introduces several design challenges, such as how to allow the operator to quickly switch between control modes and configure different types of assistance without occluding the view of the remote environment.
@@
The contribution of this work is the design and evaluation of a framework to incorporate open-world shared control into immersive robot avatars. To address the central design challenges highlighted above, we created an in-headset menu that allows the operator to launch and configure assistive actions using the same controllers they use to directly move the robot (Fig. 1). We implement assistive actions based on geometric affordances that are agnostic to object identity, allowing them to work in a wide range of scenarios. Affordances are rendered as augmented reality (AR) markers in the operator’s immersive view when the user is configuring action targets. We further enhance the fluency of this interface using an “autocomplete” predictive menu that predicts the operator’s intent in the context of the current scene and history [naughton_structured_2022]. We incorporate this framework into an avatar system and evaluate novice users on long-form tasks that require many uses of the assistive actions. Human subjects testing $(N=19)$ verifies that our approach, with and without the predictive menu, increases task success rates and system usability, and decreases task completion times and operator workload over standard direct control interfaces while preserving the operator’s self-reported sense of presence in the remote environment.
@@ II Related Work
The recent ANA Avatar XPRIZE competition spurred rapid development of teleoperated avatar robots capable of transporting basic human manipulation skills to remote environments [XPRIZESystemsPaper2023]. As the competition emphasized immersion and presence, most teams made very little or no use of shared control, instead opting to give as much direct control to the operator as possible. This choice makes the systems open-world, immersive, and intuitive, but users still struggle to perform tasks through the robot as proficiently as they would in-person [XPRIZESystemsPaper2023]. Shared control methods could hypothetically assist in operator proficiency while preserving desirable aspects of immersion, but mechanisms for achieving such integration are not well studied.
Operator assistance for non-immersive interfaces has received much attention in the literature. A significant line of work addresses reaching for an object [draganTeleoperationIntelligentCustomizable2013], especially when the operator’s interface has fewer DoFs than the robot [hauserRecognitionPredictionPlanning2013, javdani_shared_2015, quereSharedControlTemplates2020, Jeon-RSS-20]. In the avatar context, this is not normally a concern because the operator has access to high DoF input devices. Other research provides assistance for complex tasks but requires pre-programmed information about the environment and target objects [quereSharedControlTemplates2020, bustamante_cats_2022, huang_evaluation_2019]. For example, [quereSharedControlTemplates2020] presents a system that can perform complicated tasks like opening a door, but key frames of reference for specific objects are labeled by hand, and the state-machines describing transitions between different phases of the tasks are pre-specified. Our work seeks to relax this requirement and provide assistance in an open-world where the semantic identities and number of objects encountered in the environment are not known ahead of time. We achieve this by using more generic types of assistance, detecting affordances at runtime rather than hand labelling them at design-time.
The work of Pruks and Ryu [pruks_method_2022] is most similar to our system. Similar to our system, their work uses off-the-shelf methods to segment the environment into geometric primitives and allows the operator to apply customizable virtual fixtures between features detected in the environment and features from the robot. However, they use a screen-and-mouse interface to specify virtual fixtures and a separate haptic device to input low-level motion commands, requiring the operator to switch between two input devices. In contrast, our system uses a consistent input interface for both specifying virtual fixtures and providing low-level commands. Our system also provides an immersive interface via a virtual reality headset, rather than a standard screen interface. Finally, we also present a framework for incorporating predictive assistance into our system, which  [pruks_method_2022] did not consider.
@@ III Interface Design
Figure 2: System diagram showing how different interface elements control the robot. Operators use their own head and hand to control the robot’s head and hand, and use a button on their hand controller to interact with the assistive menu. The Perception Module detects affordances in the environment to display possible assistive actions to the operator. [Best viewed in color.]
@@
We build our work on the TRINA avatar system [AVATRINASystemsPaper], in which the robot is comprised of two Franka Emika Panda arms, a Robotiq 2F-140 parallel-jaw gripper, an anthropomorphic Psyonic Ability Hand, a Waypoint Vector omnidirectional wheeled base, and a custom-built three DoF neck and head assembly. A human operator controls TRINA using a virtual reality (VR) head-mounted display (HMD) that shows the view of TRINA’s environment from stereo head cameras. They control the robot’s head directly via HMD motion and use VR controllers to move the arms. The operator station is connected to the Internet via Ethernet and the robot is connected via WiFi or an Ethernet tether.
@@ III-A Direct Teleoperation (DT)
The default control mode is the direct teleoperation scheme described in [AVATRINASystemsPaper]. To simplify novice operator training, in our experiments, we only activate the robot’s right arm, parallel-jaw gripper, and head. The operator wears a VR HMD and the robot’s head tracks the operator’s head orientation. The operator uses a clutched system to control the arm: while holding down a foot pedal, the operator moves a VR controller, shown in Fig. 2, to move the robot’s hand target. This motion is computed relative to the controller’s pose when the operator first presses the pedal. A lower-level controller then attempts to reach this target. The operator can also velocity-control the parallel-jaw gripper using a joystick on the controller, pushing it right to inch the gripper closed, and left to inch it open.
The robot estimates the net force applied to its end effector to provide force feedback via two modalities: First, the controller vibrates with an intensity proportional to the estimated force magnitude (clipped between 10 and 30 N). Second, a virtual red hemisphere around the operator’s controller shows the direction of the applied force, and becomes more opaque as the magnitude of the force increases.
@@ III-B Manual Menu (MM)
Using the direct teleoperation interface alone, operators can achieve some manipulation tasks [AVATRINASystemsPaper], but complicated tasks, such as writing, are still quite difficult. To aid the operator, we created an interface to allow them to execute assistive actions. Guided by previous research [komerska_study_2004], we designed a hierarchical pie menu fixed to the operator’s head, shown in Fig. 3. By making the menu hierarchical, we minimize the number of simultaneously displayed icons to keep the operator’s view of the remote environment unobstructed. The operator interacts with the menu using a “laser pointer” emanating from their controller to point at different icons, and clicks the B button on their controller to select them. The operator can bring up this menu by clicking the B button at any time and can close it by selecting the “Close” icon. This menu design allows the operator to configure the menu using the same interface they use to provide low-level commands to the robot, eliminating any need to switch between interfaces during operation. Clicking other icons gives the operator access to different submenus.
@@ III-C Predictive Menu (PM)
While the manual menu provides access to all possible actions, it can be overwhelming and slow, especially for novice users. To alleviate this, we designed a third interface that uses an action predictor, described in section V, to predict the operator’s intent and present them with a reduced menu that only includes the four most likely actions. If the operator’s desired action is not in this set, they can still access the manual menu as a fallback. With this menu, when the operator clicks B, the top four actions are shown instead of the manual menu, as shown in 1(c) and Fig. 3. Whenever the operator hovers over an icon corresponding to an action, all other icons (and affordances) dim to 10% opacity. Selecting any icon closes the menu and sends the action to the robot which then executes it.
@@ IV Assistive Actions
We implemented three kinds of assistive actions: constrained teleoperation, snapping to planes, and snapping to circles. The use of geometric affordances to provide assistance allows the use of these actions in an open-world context, where the semantic meaning of objects in the environment is unknown. The constrained teleoperation and plane snapping actions were previously described in [naughton_structured_2022], and so are only briefly covered here.
@@
The constrained teleoperation action, teleop(sens, x, y, z, roll, pitch, yaw) accepts 7 Boolean parameters modifying the operator’s direct control of the arm. During this action, the operator controls the gripper’s target pose by moving a VR controller with their own arm. When the sens parameter is true, the arm’s end-effector motion is isotropically scaled to $0.25$ of the operator’s input motion to enable precise manipulation. The remaining parameters toggle constraints on the end-effector motion, activating guidance virtual fixtures to simplify operation [bowyerActiveConstraintsVirtual2014].
@@
The plane snapping action, snap_to_plane(p) accepts a plane detected from a point-cloud of the environment by a clustering method [fengFastPlaneExtraction2014]. This point-cloud is sensed by the “affordance camera” shown in Fig. 2, an Intel RealSense L515 mounted below the robot’s neck, pointed at the center of the robot’s workspace. The plane extraction algorithm updates the set of detected planes once every 5 seconds. This action aligns the forward direction of the gripper with the normal of the detected plane and moves it so that its tool tip is $d_{s}$ m away from the plane to prepare the operator to perform manipulation on or near the plane’s surface. For the tasks considered here we found $d_{s}=0.15$ m to work well. Fig. 4 illustrates this process in 2D. The robot uses a sampling-based planner to find a path to reach this target or reports that no path was found after 10 s.
@@
Lastly, the snap_to_circle(c) action accepts a circle detected from the environment, aligns the gripper’s forward direction with the circle’s axis, and centers the gripper on the circle to prepare the operator to perform rotating manipulations about the circle’s axis. Our system detects circles from RGBD images from the affordance camera once every 5 seconds. The system segments the RGB image using the Segment Anything Model (SAM) [kirillov2023segany] and converts the RGBD image into a point-cloud. For each image mask, the corresponding points are selected, and the plane supported by the most points is found. The inliers of this plane are computed as the points in the mask within $d_{\text{in}}=5\text{ mm}$ of the plane and projected to the plane. The convex hull of these projected points is found and the circle is discarded if this hull’s “circularity” ($\frac{4\pi\cdot\text{Area}}{\text{Perimeter}^{2}}$ [opencv_library]) is below $c_{\text{min}}=0.9$. The minimum enclosing circle of the hull is computed and circles with radii greater than $r_{\text{max}}=7\text{ cm}$ are discarded. To remove duplicates, this candidate circle is compared against previously detected circles. Circles are considered similar if the masks from which they were detected overlap, their centers are within $\Delta_{\text{c}}=5\text{ cm}$, and their radii are within $\Delta_{\text{rad}}=1\text{ cm}$. Among similar circles, the one with the largest ratio of inliers to points in the mask is kept. Once a circle has been selected, the robot computes a target end-effector pose in the same manner as the snap_to_plane action, additionally moving the target so that the projection of the tool tip to the plane of the circle coincides with the circle’s center. Fig. 4 demonstrates this action in 2D.
@@
Figure 4: 2D illustration of the snap_to_plane and snap_to_circle actions. Both align TRINA’s gripper with the normal of the selected affordance, but snap_to_circle centers the gripper on the circle while snap_to_plane only moves it closer to the plane. Here we used $d_{s}=0.15$ m. [Best viewed in color.]
@@ V Intent Prediction
To populate the predictive menu, we require an action predictor that can predict multiple likely actions. Additionally, since the set of affordances is not known until runtime, the predictor must be open-world, i.e. able to predict over an open set of objects. We employ the structured prediction method of [naughton_structured_2022] as it was found to have strong performance in open-world scenarios on similar tasks.
Actions are defined by a type and a collection of parameters, $\overline{\psi}$, which may be different for each action type. We limit the set of $n$ types a priori and dynamically detect the set of feasible parameters for each type, corresponding to detected affordances. To predict an action given the robot’s current context vector, $x$, the method uses $n$ parameter scoring neural networks, $\{G^{(i)}(x,\overline{\psi})\}_{i=1}^{n}$, and an action network, $A(x)$. $A(x)$ produces an $n$-dimensional output vector with each element representing the overall score for an action type. Each $G^{(i)}(x,\overline{\psi})$ predicts a scalar score for parameter collections of a particular action type. To score a complete action, the appropriate scores are summed, $s=e_{i}^{\intercal}A(x)+G^{(i)}(x,\overline{\psi})$, where $e_{i}$ is the $i$th standard basis vector.
To train and evaluate our predictor, three expert operators (paper authors) collected a dataset of 150 action sequences across three different tasks: unscrewing a jar lid, writing “IML” on a whiteboard, and plugging a cord into an electrical socket. Each sequence was collected in a highly cluttered environment that contained many different distractor objects with varied compositions and arrangements. The specific target objects used were also modified (for example, varying which jars were used). The scoring function was trained using a maximum margin loss function to output high scores for actions observed in the demonstrations [naughton_structured_2022].